# Evaluating the Posterior Sampling Ability of Plug&Play Diffusion Methods in Sparse-View CT

### v1

@@ 4 Posterior evaluation criteria
$p$ PSNR $\uparrow$ NMC $\xrightarrow{}1$ FID $\downarrow$ CMMD $\downarrow$ PSNR $\uparrow$ NMC $\xrightarrow{}1$ FID $\downarrow$ CMMD $\downarrow$ PSNR $\uparrow$ NMC $\xrightarrow{}1$ FID $\downarrow$ CMMD $\downarrow$
LDCT $180$ $\mathbf{37.70}$ $\mathbf{0.99}$ $\mathbf{0.57}$ $\mathbf{0.002}$ $33.09$ $1.11$ $0.58$ $0.003$- -- -- --
$90$ $\mathbf{36.57}$ $\mathbf{1.00}$ $1.06$ $0.088$ $33.05$ $1.11$ $\mathbf{0.57}$ $\mathbf{0.003}$- -- -- --
$30$ $\mathbf{34.21}$ $\mathbf{1.06}$ $2.27$ $0.386$ $32.89$ $1.12$ $\mathbf{0.54}$ $\mathbf{0.001}$- -- -- --
$18$ $\mathbf{33.10}$ $\mathbf{1.09}$ $2.61$ $0.449$ $32.66$ $1.12$ $\mathbf{0.55}$ $0.004$ $29.67$ $1.92$ $\mathbf{0.55}$ $\mathbf{0.0012}$
$12$ $32.19$ $1.15$ $3.90$ $0.611$ $\mathbf{32.37}$ $\mathbf{1.11}$ $0.61$ $0.012$ $29.36$ $2.01$ $\mathbf{0.54}$ $\mathbf{0.0003}$
$6$ $30.62$ $1.32$ $5.94$ $0.733$ $\mathbf{31.61}$ $\mathbf{1.14}$ $0.82$ $0.047$ $28.82$ $2.19$ $\mathbf{0.54}$ $\mathbf{0.0003}$
$3$ $28.15$ $1.53$ $7.66$ $0.829$ $\mathbf{30.23}$ $\mathbf{1.22}$ $1.40$ $0.125$ $27.75$ $3.02$ $\mathbf{0.55}$ $\mathbf{0.0002}$
$1$ $20.13$ $1.72$ $8.54$ $0.754$ $\mathbf{23.32}$ $\mathbf{1.53}$ $6.34$ $0.563$ $20.27$ $14.89$ $\mathbf{1.00}$ $\mathbf{0.0070}$
LIDC $180$ $\mathbf{35.79}$ $\mathbf{1.06}$ $1.34$ $0.022$ $29.06$ $1.50$ $\mathbf{0.87}$ $\mathbf{0.012}$- -- -- --
$90$ $\mathbf{34.44}$ $\mathbf{1.09}$ $1.00$ $0.009$ $29.01$ $1.51$ $\mathbf{0.86}$ $\mathbf{0.012}$- -- -- --
$30$ $\mathbf{31.65}$ $\mathbf{1.20}$ $2.35$ $0.166$ $28.83$ $1.50$ $\mathbf{0.84}$ $\mathbf{0.010}$- -- -- --
$18$ $\mathbf{30.19}$ $\mathbf{1.29}$ $3.25$ $0.278$ $28.58$ $1.46$ $0.89$ $0.010$ $26.02$ $3.01$ $\mathbf{0.54}$ $\mathbf{0.0007}$
$12$ $\mathbf{28.80}$ $\mathbf{1.42}$ $4.44$ $0.360$ $27.96$ $\mathbf{1.42}$ $1.01$ $0.012$ $25.39$ $3.79$ $\mathbf{0.53}$ $\mathbf{0.0005}$
$6$ $\mathbf{26.53}$ $1.65$ $5.67$ $0.411$ $26.50$ $\mathbf{1.36}$ $1.20$ $0.019$ $24.26$ $4.71$ $\mathbf{0.53}$ $\mathbf{0.0005}$
$3$ $22.83$ $1.83$ $5.97$ $0.421$ $\mathbf{23.61}$ $\mathbf{1.41}$ $1.39$ $0.031$ $22.15$ $7.12$ $\mathbf{0.56}$ $\mathbf{0.0008}$
$1$ $17.22$ $0.29$ $4.90$ $0.337$ $\mathbf{17.89}$ $\mathbf{1.62}$ $2.80$ $0.164$ $17.02$ $14.58$ $\mathbf{0.67}$ $\mathbf{0.0019}$
JET $180$ $\mathbf{37.66}$ $\mathbf{1.11}$ $4.87$ $0.151$ $30.39$ $1.90$ $\mathbf{1.03}$ $\mathbf{0.040}$- -- -- --
$90$ $\mathbf{35.53}$ $\mathbf{1.15}$ $4.86$ $0.086$ $30.37$ $1.90$ $\mathbf{1.01}$ $\mathbf{0.040}$- -- -- --
$30$ $\mathbf{33.41}$ $\mathbf{1.21}$ $2.29$ $0.090$ $30.28$ $1.90$ $\mathbf{0.96}$ $\mathbf{0.042}$- -- -- --
$18$ $\mathbf{31.68}$ $\mathbf{1.27}$ $5.06$ $0.542$ $29.92$ $1.84$ $1.02$ $0.044$ $26.31$ $2.11$ $\mathbf{0.62}$ $\mathbf{0.008}$
$12$ $\mathbf{30.20}$ $\mathbf{1.40}$ $10.23$ $1.096$ $29.26$ $1.77$ $1.28$ $0.060$ $25.82$ $5.25$ $\mathbf{0.64}$ $\mathbf{0.009}$
$6$ $\mathbf{27.66}$ $1.95$ $14.20$ $1.733$ $27.58$ $\mathbf{1.64}$ $2.23$ $0.141$ $24.99$ $6.57$ $\mathbf{0.63}$ $\mathbf{0.009}$
$3$ $25.66$ $2.42$ $14.21$ $2.298$ $\mathbf{25.86}$ $\mathbf{1.59}$ $3.80$ $0.235$ $24.21$ $7.99$ $\mathbf{0.58}$ $\mathbf{0.007}$
$1$ $23.85$ $2.44$ $13.96$ $1.941$ $\mathbf{24.13}$ $\mathbf{1.84}$ $12.89$ $1.061$ $23.33$ $9.72$ $\mathbf{0.56}$ $\mathbf{0.005}$
Table 2: Posterior evaluation of three state-of-the-art PnP diffusion methods [10, 11, 12] in SVCT. See Sec. 5.4 for details.
5 Experiments
@@ 5.1 Training Datasets
$\sigma_{y}=\frac{1}{100\times I}\sum_{i\in\mathcal{D}}\left(\text{max}(\mathbf{H}_{180}\mathbf{x}_{i})-\text{min}(\mathbf{H}_{180}\mathbf{x}_{i})\right).$ (8)
@@ 5.3 Technical details on PnP diffusion methods
We consider three state-of-the-art PnP diffusion methods: MCG [10], DPS [11] and $\Pi$G [12]. They essentially differ in the respective approximations of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ (see Tab. 1). MCG and DPS have handcrafted weights $\alpha(\mathbf{x}_{t},\mathbf{y}_{p})$ originally defined as $\alpha_{\text{MCG}}(\mathbf{x}_{t},\mathbf{y}_{p})=0.1/\|\mathbf{H}_{p}^{{\dagger}}(\mathbf{y}_{p}-\mathbf{H}_{p}\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}$ and $\alpha_{\text{DPS}}(\mathbf{x}_{t},\mathbf{y}_{p})=1/\|\mathbf{y}_{p}-\mathbf{H}_{p}\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})\|_{2}$. We found these values to be non-optimal and fine-tuned them on each dataset for each value of $p$. $\Pi$G proposes time-dependent weights: $r_{t}=\sqrt{\sigma_{t}^{2}/\sigma_{t}^{2}+1}$ with $\sigma_{t}$ the diffusion noise, we did not modify this scheme. To fairly compare the three methods, in all the experiments, we use the same accelerated ancestral sampler[43, 12] with 100 noise scales with $\sigma_{min}=0.01$ and $\sigma_{max}=1348$.
@@ 5.4 Quantitative evaluation
We aim at evaluating the approximate posterior of each PnP diffusion method [10, 11, 12] as a function of the number of projections $p$ (in the experiments $p=1,3,6,12,18,30,90,180$). To do so, for each dataset, we proceed as follows. We unconditionally sample a set $\mathcal{X}$ of $50$k images using the time-dependent prior score network $\mathbf{s}_{\boldsymbol{\theta}}$, as explained in Sec. 4. From $\mathcal{X}$, for each value of $p$, we compute a set of $50$k $p$-projections sinograms, $\mathcal{Y}_{p}$, using the SVCT measurement model Eq. (1). Then for each method $m\in$ [10, 11, 12], we compute a set of $50$k posterior samples, $\overset{\sim}{\mathcal{X}}_{p,m}$, using sinograms of $\mathcal{Y}_{p}$ as conditions. To evaluate Eq. (6) we compute the FID [29] and the CMMD [36] between $\mathcal{X}$ and $\overset{\sim}{\mathcal{X}}_{p,m}$. To evaluate Eq. (7), we compute the NMC (see Sec. 4) between $\mathcal{Y}_{p}$ and $\overset{\sim}{\mathcal{X}}_{p,m}$. We also compute the average PSNR between $\mathcal{X}$ and $\overset{\sim}{\mathcal{X}}_{p,m}$. Due to the limit of space, we did not include the SSIM, as it follows the same behaviour as the PSNR. The evaluation results are summarized in Tab. 2. Let us highlight that to compute $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$, $\Pi$G needs to solve a linear system whose size increases with $p$ (see Tab. 1), thus we could only compute the metrics for $p=\{1,3,6,12,18\}$. In theory MCG needs to perform a similar operation (pseudo-inverse), but in practice, as suggested in [10], we use a Filtered Back-Projection [14] (FBP), that significantly speeds-up the sampling.
One can see the average PSNR of each method is monotonically increasing with $p$. This is expected since $p$ controls the amount of information contained in sinograms.
Results of MCG - The NMC is close to 1 which indicates conditioning sinograms were correctly taken into account during sampling. However, the FID and the CMMD significantly increase when $p$ decreases. This dependency on $p$ is unexpected and shows that MCG’s approximate posterior deviates from the true posterior when $p$ decreases.
Results of DPS - The NMC is close to 1 which indicates conditioning sinograms were correctly taken into account during sampling. The FID and the CMMD are low and only slightly increase when $p$ decreases (except for $p=1$). These findings tend to show DPS’s posterior is close to the true posterior, even for a low number of projections (except in the extreme case $p=1$).

### v2

@@ 4 Posterior evaluation
$p$ $\text{PLS}_{\text{NMC}}$ $\text{PPS}_{\text{FID}}$ $\text{PPS}_{\text{CMMD}}$ $\text{PLS}_{\text{NMC}}$ $\text{PPS}_{\text{FID}}$ $\text{PPS}_{\text{CMMD}}$ $\text{PLS}_{\text{NMC}}$ $\text{PPS}_{\text{FID}}$ $\text{PPS}_{\text{CMMD}}$
$\xrightarrow{}1$downarrow $\downarrow$ $\xrightarrow{}1$downarrow $\downarrow$ $\xrightarrow{}1$downarrow $\downarrow$
LDCT $180$ $\mathbf{0.89}$ $\mathbf{0.57}$ $\mathbf{0.002}$ $1.14$ $0.58$ $0.003$- -- --
$90$ $0.82$ $1.06$ $0.088$ $\mathbf{1.12}$ $\mathbf{0.57}$ $\mathbf{0.003}$- -- --
$30$ $0.65$ $2.27$ $0.386$ $\mathbf{1.05}$ $\mathbf{0.54}$ $\mathbf{0.001}$- -- --
$18$ $0.45$ $2.61$ $0.449$ $\mathbf{0.98}$ $\mathbf{0.55}$ $0.004$ $1.92$ $\mathbf{0.55}$ $\mathbf{0.0012}$
$12$ $0.32$ $3.90$ $0.611$ $\mathbf{0.90}$ $0.61$ $0.012$ $2.01$ $\mathbf{0.54}$ $\mathbf{0.0003}$
$6$ $0.20$ $5.94$ $0.733$ $\mathbf{0.72}$ $0.82$ $0.047$ $2.19$ $\mathbf{0.54}$ $\mathbf{0.0003}$
$3$ $0.15$ $7.66$ $0.829$ $\mathbf{0.55}$ $1.40$ $0.125$ $3.02$ $\mathbf{0.55}$ $\mathbf{0.0002}$
$1$ $0.20$ $8.54$ $0.754$ $\mathbf{0.24}$ $6.34$ $0.563$ $14.89$ $\mathbf{1.00}$ $\mathbf{0.0070}$
LIDC $180$ $\mathbf{0.87}$ $1.34$ $0.022$ $1.48$ $\mathbf{0.87}$ $\mathbf{0.012}$- -- --
$90$ $\mathbf{0.80}$ $1.00$ $0.009$ $1.45$ $\mathbf{0.86}$ $\mathbf{0.012}$- -- --
$30$ $0.63$ $2.35$ $0.166$ $\mathbf{1.35}$ $\mathbf{0.84}$ $\mathbf{0.010}$- -- --
$18$ $0.46$ $3.25$ $0.278$ $\mathbf{1.20}$ $0.89$ $0.010$ $3.01$ $\mathbf{0.54}$ $\mathbf{0.0007}$
$12$ $0.34$ $4.44$ $0.360$ $\mathbf{1.04}$ $1.01$ $0.012$ $3.79$ $\mathbf{0.53}$ $\mathbf{0.0005}$
$6$ $0.22$ $5.67$ $0.411$ $\mathbf{0.74}$ $1.20$ $0.019$ $4.71$ $\mathbf{0.53}$ $\mathbf{0.0005}$
$3$ $0.15$ $5.97$ $0.421$ $\mathbf{0.51}$ $1.39$ $0.031$ $7.12$ $\mathbf{0.56}$ $\mathbf{0.0008}$
$1$ $0.21$ $4.90$ $0.337$ $\mathbf{0.27}$ $2.80$ $0.164$ $14.58$ $\mathbf{0.67}$ $\mathbf{0.0019}$
JET $180$ $\mathbf{0.98}$ $4.87$ $0.151$ $1.70$ $\mathbf{1.03}$ $\mathbf{0.040}$- -- --
$90$ $\mathbf{0.96}$ $4.86$ $0.086$ $1.68$ $\mathbf{1.01}$ $\mathbf{0.040}$- -- --
$30$ $\mathbf{0.90}$ $2.29$ $0.090$ $1.63$ $\mathbf{0.96}$ $\mathbf{0.042}$- -- --
$18$ $\mathbf{0.81}$ $5.06$ $0.542$ $1.51$ $1.02$ $0.044$ $2.11$ $\mathbf{0.62}$ $\mathbf{0.008}$
$12$ $0.63$ $10.23$ $1.096$ $\mathbf{1.36}$ $1.28$ $0.060$ $5.25$ $\mathbf{0.64}$ $\mathbf{0.009}$
$6$ $0.30$ $14.20$ $1.733$ $\mathbf{1.07}$ $2.23$ $0.141$ $6.57$ $\mathbf{0.63}$ $\mathbf{0.009}$
$3$ $0.17$ $14.21$ $2.298$ $\mathbf{0.83}$ $3.80$ $0.235$ $7.99$ $\mathbf{0.58}$ $\mathbf{0.007}$
$1$ $0.21$ $13.96$ $1.941$ $\mathbf{0.48}$ $12.89$ $1.061$ $9.72$ $\mathbf{0.56}$ $\mathbf{0.005}$
Table 2: Posterior evaluation of three state-of-the-art PnP diffusion methods [10, 11, 12] in SVCT. See Sec. 4.4 for details. Corrected $\text{PLS}_{\text{NMC}}$ results following the identification of a computational bug in the official publication. The $\text{PLS}_{\text{NMC}}$ corrected values presented here replace those in the original table, which contained errors due to the bug. These corrected results do not affect the overall conclusions of the study.
@@ 4.1 Training Datasets
$\sigma_{y}=\frac{1}{100\times{N}}\sum_{i\in\mathcal{D}}\left(\text{max}(\mathbf{H}_{180}\mathbf{x}_{i})-\text{min}(\mathbf{H}_{180}\mathbf{x}_{i})\right).$ (7)
@@ 4.3 Technical details on PnP diffusion methods
We consider three SOTA PnP diffusion methods: MCG [10], DPS [11] and $\Pi$[12]. They essentially differ in the respective approximations of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$ (see Tab. 1). MCG and DPS have handcrafted weights originally defined as $\alpha_{\text{MCG}}(\mathbf{x}_{t},\mathbf{y}_{p})=0.1/\|\mathbf{H}_{p}^{{\dagger}}(\mathbf{y}_{p}-\mathbf{H}_{p}\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}$ and $\alpha_{\text{DPS}}(\mathbf{x}_{t},\mathbf{y}_{p})=1/\|\mathbf{y}_{p}-\mathbf{H}_{p}\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})\|_{2}$. We found these values to be non-optimal and fine-tuned them on each dataset for each value of $p$. $\Pi$G proposes time-dependent weights: $r_{t}=\sqrt{\sigma_{t}^{2}/\sigma_{t}^{2}+1}$ with $\sigma_{t}$ the diffusion noise, we did not modify this scheme. To fairly compare the three methods, in all the experiments, we use the same accelerated ancestral sampler[43, 12] with 100 noise scales with $\sigma_{min}=0.01$ and $\sigma_{max}=1348$.
@@ 4.4 Quantitative evaluation
We aim at evaluating the approximate posterior of each PnP diffusion method [10, 11, 12] as a function of the number of projections $p$ (in the experiments $p=1,3,6,12,18,30,90,180$). To do so, for each dataset, we proceed as follows. We unconditionally sample a set $\mathcal{X}$ of $50$k images using the time-dependent prior score network $\mathbf{s}_{\boldsymbol{\theta}}$, as explained in Sec. 4. From $\mathcal{X}$, for each value of $p$, we compute a set of $50$k $p$-projections sinograms, $\mathcal{Y}_{p}$, using the SVCT measurement model Eq. (1). Then for each method $m\in$ [10, 11, 12], we compute a set of $50$k posterior samples, $\overset{\sim}{\mathcal{X}}_{p,m}$, using sinograms of $\mathcal{Y}_{p}$ as conditions. To evaluate Eq. (5) we compute the FID [29] and the CMMD [36] between $\mathcal{X}$ and $\overset{\sim}{\mathcal{X}}_{p,m}$. To evaluate Eq. (6), we compute the NMC (see Sec. 4) between $\mathcal{Y}_{p}$ and $\overset{\sim}{\mathcal{X}}_{p,m}$. The evaluation results are summarized in Tab. 2. Let us highlight that to compute $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{p}|\mathbf{x}_{t})$, $\Pi$G needs to solve a linear system whose size increases with $p$ (see Tab. 1), thus we could only compute the metrics for $p=\{1,3,6,12,18\}$. In theory MCG needs to perform a similar operation (pseudo-inverse) but, as suggested in [10], we use a Filtered Back-Projection [14] (FBP), that significantly speeds-up the sampling.
Results of MCG - At $p=180$, the NMC is slightly below $1$, indicating a mild overfitting regime. As $p$ decreases, MCG shows increasing overfitting, with the NMC approaching 00. Both the FID and CMMD metrics rise significantly as $p$ decreases, which aligns with the NMC results. Indeed, greater overfitting leads to more correlated noise in the posterior samples. When this correlated noise is projected back into sinogram space, it manifests as the overfitted noise and hence differs from prior samples. However, the observed dependency on $p$ is unexpected, indicating that MCG’s approximate posterior begins to diverge from the true posterior as $p$ decreases.
Results of DPS - At $p=180$, the NMC is slightly over $1$, indicating a mild underfitting regime. As $p$ decreases, DPS transitions into overfitting regime.Both the FID and CMMD metrics increase significantly in the overfitting scenario. This is again due to the overfitting correlated noise in posterior samples. The NMC and FID, CMMD seems to be ideal around $p=18$ but we still observe a dependency in $p$.